4.) Two parallel wires are separated by a distance of 0.05 m. Wire 1 carries a carries current of 12.0 A. Wire 2 carries a current of 6.0 A flowing in the opposite direction as shown in the illustration. Find the (a) magnitude and direction of the magnetic field created by wire 1 at wire 2, (b) force exerted by wire 1 on the 2.0 m long section of wire 2, and (c) force exerted by wire 2 on wire 1.



5.) What should be the length of an ideal solenoid consisting of 1250 turns for the magnetic field in its interior to be 0.025 T? A current of 2.0 A is flowing through the solenoid.

1
Expert's answer
2022-04-11T12:04:53-0400

4.

a)

"B_1=\frac{\mu_0I_1}{2\pi d}=48~\mu T,"

b)

"F_{12}=\frac{\mu_0I_1I_2 l}{2\pi d}=0.58~mA,"

c)

"F_{21}=\frac{\mu_0I_2I_1 l}{2\pi d}=0.58~mA."


5.

"B=\frac{\mu_0 IN}l,\implies"

"l=\frac{\mu_0 IN}B=126~mm."